OpenStudy (kweb2017):

A 22 gram sample of a substance that's used to sterilize surgical instruments has a k-value of 0.1363. Formula: N=N0e^-kt Find the substance's half-life, in days. Round to nearest tenth. N0=initial mass (at time t=0) N=mass at time t k=a positive constant that depends on the substance itself and on the units used to measure time t=time, in days

satellite73 (satellite73):

set \[\large e^{-0.1363}=0.5\] and solve for \(t\)

OpenStudy (kweb2017):

how...

satellite73 (satellite73):

half life means the time it takes to get to one half or \(0.5\)

satellite73 (satellite73):

takes two steps only

satellite73 (satellite73):

first write in equivalent logarthmic form as \[-0.1363t=\ln(0.5)\]

satellite73 (satellite73):

the to solve for \(t\) divide to get \[\large t=\frac{\ln(0.5)}{-0.1363}\]

satellite73 (satellite73):

and of course a calculator

OpenStudy (kweb2017):

can you walk me through all of that? I'm sorry I'm really struggling with this section in math..

satellite73 (satellite73):

ok although i already wrote the steps, i will add some explanation, maybe that will help

OpenStudy (kweb2017):

yes please

satellite73 (satellite73):

the inverses of the function \(f(x)=e^x\) is the natural log function \(\ln(x)\)

satellite73 (satellite73):

one thing you should be able to do is to switch from exponential form to logarithmic form quickly and easily for example \[10^3=1000\] same as \[\log_{10}(1000)=3\]

satellite73 (satellite73):

\[2^4=16\\ \log_2(16)=4\] they same the same thing

OpenStudy (kweb2017):

would the answer be -80.70?

satellite73 (satellite73):

no the answer is not negative

OpenStudy (kweb2017):

so its 80.70

satellite73 (satellite73):

\[e^x=y\\ \ln(y)=x\]

satellite73 (satellite73):

\[e^{-0.1363t}=0.5\\ -0.1363t=\ln(.5)\]

OpenStudy (kweb2017):

hold on hold on what do i need to plug in this equation t=ln(0.5)/0.1363

satellite73 (satellite73):

yes

satellite73 (satellite73):

oh i see the mistake

OpenStudy (kweb2017):

what do i need to plug in?

satellite73 (satellite73):

the denominator should be negative

OpenStudy (kweb2017):

what?

satellite73 (satellite73):

lets go back to the start \[e^{-0.136t}=0.5\\ -0.1363t=\ln(0.5)\\ t=\frac{\ln(0.5)}{-0.1363}\]

OpenStudy (kweb2017):

ok ok what do i plug into this equation?

satellite73 (satellite73):

you have nothing to "plug in" type in to your calculator (not sure which one you are using ) \[\ln(.5)\div (-.1363)\]

satellite73 (satellite73):

i use this http://www.wolframalpha.com/input/?i=ln%28.5%29%2F-.1363

OpenStudy (kweb2017):

what is ln

OpenStudy (kweb2017):

5.09?

satellite73 (satellite73):

the natural log function , log base \(e\) \[\log_e(x)=\ln(x)\]

OpenStudy (kweb2017):

oooooh ok

OpenStudy (kweb2017):

so my answer is 5.08?

satellite73 (satellite73):

that is what i got, yes

OpenStudy (kweb2017):

final answer?

satellite73 (satellite73):

or correctly rounded \(5.09\)

OpenStudy (kweb2017):

ok thank you

satellite73 (satellite73):

oh wait

satellite73 (satellite73):

it asks "to the nearest tenth" so \(5.1\)

OpenStudy (kweb2017):

it said it was incorrect

OpenStudy (kweb2017):

oh man lol

OpenStudy (kweb2017):

oh well it'll buff thank you for teaching me. X

satellite73 (satellite73):

now i have a question

OpenStudy (kweb2017):

ok :)

satellite73 (satellite73):

what is your instagram?

OpenStudy (kweb2017):

lol give me yours

satellite73 (satellite73):

guess

OpenStudy (kweb2017):

satellite73?

satellite73 (satellite73):

bingo